OpenStudy (anonymous):

Choose the equation of the line passing through the point (-1, -2) and perpendicular to y =1/4 x + 2.

OpenStudy (anonymous):

well y=mx+b is your slope. and you are now given the x and y values so just plug it in

OpenStudy (anonymous):

y = 4x - 6 y = -4x - 6 y = -4x + 2 y = 4x thats the multiple choices

OpenStudy (anonymous):

I'm a bit confused sorry :l

OpenStudy (anonymous):

you dont understand it?

OpenStudy (anonymous):

It;s not that, I just don't know how I would change it

OpenStudy (anonymous):

me 2

OpenStudy (anonymous):

Well I can tell you that the m is the original and the b is the changing factor so try to plug that in and maybe when you graph it you can get the right answer but my guess would be c

OpenStudy (anonymous):

ok thxx

OpenStudy (anonymous):

Sure thing

OpenStudy (anonymous):

so the solution will be y-(-2)=x-(-1)

OpenStudy (anonymous):

- 1/4 both side?

OpenStudy (anonymous):

im sorry i dive

OpenStudy (anonymous):

divide

OpenStudy (anonymous):

can u do it for for me koz i really dont understand